OpenStudy (anonymous):

Thomas walks his dog for 15 minutes today- 5 more minutes than yesterday. Express as a percent his walking time today compared to his walking time yesterday. Round to the nearest percent. A) 125% B) 200% C) 133% D) 150% (I understand how to do this, but apparently I am using the wrong numbers and get something way under these percents.)

OpenStudy (anonymous):

I think A or B not sure witch one tho

OpenStudy (mathmath333):

today - 15 min yesterday - 10 min % age walking time today compared to his walking time yesterday \(\large\tt \begin{align} \color{black}{ today=yesterday\times \dfrac{x}{100}\%\\ 15=10\times \dfrac{x}{100}\%\\ x\%=\dfrac{1500}{10}\\ x\%=150\\ }\end{align}\)